OpenStudy (anonymous):

Write an equation for the following translation of y=-|x| (2 units down)

OpenStudy (solomonzelman):

Rules of \(\large\color{black}{ \rm shifts }\) from \(\large\color{black}{ \rm f(x) }\) to \(\large\color{black}{ \rm g(x) }\). \(\large\color{black}{ \rm f(x)=-\left| x \right| ~~~~~~~~~\rm{\Longrightarrow}~~~~~~~~\rm g(x)=-\left| x \color{blue}{ -~\rm{c} }\right| }\) \(\large\color{blue}{ ~\rm {c} }\) units to the \(\normalsize\color{blue}{ \rm right }\). \(\large\color{black}{ \rm f(x)=-\left| x \right| ~~~~~~~~~\rm{\Longrightarrow}~~~~~~~~\rm g(x)=-\left| x \color{blue}{ +~\rm{c} }\right| }\) \(\large\color{blue}{ ~\rm {c} }\) units to the \(\normalsize\color{blue}{ \rm left }\). \(\large\color{black}{ \rm f(x)=-\left| x \right| ~~~~~~~~~\rm{\Longrightarrow}~~~~~~~~\rm g(x)=-\left| x \right| \color{blue}{ +~\rm{c} }}\) \(\large\color{blue}{ ~\rm {c} }\) units \(\normalsize\color{blue}{ \rm up }\). \(\large\color{black}{ \rm f(x)=-\left| x \right| ~~~~~~~~~\rm{\Longrightarrow}~~~~~~~~\rm g(x)=-\left| x \right| \color{blue}{ -~\rm{c} }}\) \(\large\color{blue}{ ~\rm{c} }\) units \(\normalsize\color{blue}{ \rm down }\).

OpenStudy (solomonzelman):

tell me if you need more help.